OpenStudy (anonymous):

girl who hates word problems can someone help me? shauntee has some money she wants to put in a 6-month certificatee of deposit (CD) account that pays 5% annual interest. she wants to earn $40 in interest to buy a telephone for her room. how much money should shauntee put in the CD account?

OpenStudy (anonymous):

40 = p * .05 * 1/2 =>p = 40 / .025 = 1,600

OpenStudy (anonymous):

can u expand on how u got the answer?

OpenStudy (experimentx):

Interest = Rate x Time x Principal

OpenStudy (experimentx):

And change Rate in to decimal

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

wait what does the 1/2 mean? that chloropyhll wrote?

OpenStudy (experimentx):

that time .. 1/2 is half year.

OpenStudy (anonymous):

hmmm thats what i was thinking put wasnt sure... :)

OpenStudy (anonymous):

wait now im confused..... when chlorophyll wrote p=40/.025=1600 does this mean that 40 was in the numerator??? if so why?

OpenStudy (anonymous):

40 = p * 5/100 * 1/2 =>40 = 5 p/ 200 => p = 40 * 200/ 5 = 1,600

OpenStudy (anonymous):

Is it easier to understand?

OpenStudy (anonymous):

ummmmmmmmm.........ok i get confused in the 2nd step

OpenStudy (anonymous):

5/100 * 1/2 = 5/200

OpenStudy (anonymous):

question......where is 100 and 200 coming 4rm????????

OpenStudy (anonymous):

money you want (40) = money you need (p) * interest (0,5) * time (1/2 year) 40 = p * 0,5 * 1/2 40 = p * 0,025 40/0,025 = p 1600 = p Money you need = 1600

OpenStudy (anonymous):

5/ 100 is from 5%

OpenStudy (anonymous):

i think $800